OpenStudy (anonymous):

Please Please help! Find an equation of the line through (2,5) and parallel to y=4x-4. write the equation using the function notation.

OpenStudy (anonymous):

I really need some help with this

hero (hero):

y = mx + b y = 4x + b 5 = 4(2) + b 5 = 8 + b 5 - 8 = b -3 = b So the parallel line is y = 4x - 3 Re-written in function notation: y = f(x), therefore f(x) = 4x - 3